Draw all possible samples of size 2 from samples  10   20  30   40   50 with replacement .

Prove that mean of samples mean is same as the population mean and form empirical relation between samples variance and population variance.


1
Expert's answer
2021-09-24T06:39:05-0400

In total there are "5^2=25" samples or pairs which are possible

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:}
 Sample & Mean & & Sample & Mean & & \\ \hline
 (10,10) & 10 & & (10,20) & 15 \\
 \hdashline
 (10,30) & 20 & & (10,40) & 25 \\
 \hdashline
 (10,50) & 30 & & (20,10) & 15 \\
 \hdashline
 (20,20) & 20 & & (20,30) & 25 \\
 \hdashline
 (20,40) & 30 & & (20,50) & 35 \\
 \hdashline
 (30,10) & 20 & & (30,20) & 25 \\
 \hdashline
 (30,30) & 30 & & (30,40) & 35 \\
 \hdashline
 (30,50) & 40 & & (40,10) & 25 \\
 \hdashline
 (40,20) & 30 & & (40,30) & 35 \\
 \hdashline
 (40,40) & 40 & & (40,50) & 45 \\
 \hdashline
 (50,10) & 30 & & (50,20) & 35 \\
 \hdashline
 (50,30) & 40 & & (50,40) & 45 \\
 \hdashline
 (50,50) & 50 & & \\
 
\end{array}"

The table shows that there are nine possible values of the sample mean "\bar{X}."


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 \bar{x} & p(\bar{x}) \\ \hline
 10 & 1/25 \\
 15 & 2/25 \\
 20 & 3/25 \\
 25 & 4/25 \\
 30 & 5/25 \\
 35 & 4/25 \\
 40 & 3/25 \\
 45 & 2/25 \\
 50 & 1/25 \\
\end{array}"

"\mu_{\bar{X}}=\sum_i\bar{x}_ip(\bar{x}_i)=10(\dfrac{1}{25})+15(\dfrac{2}{25})+20(\dfrac{3}{25})"

"+25(\dfrac{4}{25})+30(\dfrac{5}{25})+35(\dfrac{4}{25})+40(\dfrac{3}{25})+45(\dfrac{2}{25})"

"+50(\dfrac{1}{25})=30"

"\sum_i\bar{x}_i^2p(\bar{x}_i)=(10)^2(\dfrac{1}{25})+(15)^2(\dfrac{2}{25})+(20)^2(\dfrac{3}{25})"

"+(25)^2(\dfrac{4}{25})+(30)^2(\dfrac{5}{25})+(35)^2(\dfrac{4}{25})+(40)^2(\dfrac{3}{25})"

"+(45)^2(\dfrac{2}{25})+(50)^2(\dfrac{1}{25})=1000"

"Var(\bar{X})=\sigma_{\bar{X}}^2=\sum_i\bar{x}_i^2p(\bar{x}_i)-(\sum_i\bar{x}_ip(\bar{x}_i))^2"

"=1000-(30)^2=100"

"\sigma_{\bar{X}}=\sqrt{\sigma_{\bar{X}}^2}=\sqrt{100}=10"

"\mu=\dfrac{10+20+30+40+50}{5}=30"

In sampling with replacement the mean of all sample means equals the mean of the population:

"\mu_{\bar{X}}=30=\mu"

"\sigma^2=\dfrac{1}{5}((10-30)^2+(20-30)^2+(30-30)^2"

"+(40-30)^2+(50-30)^2)=200"

When sampling with replacement the variation of all sample means equals the variation of the population divided by the sample size 


"\sigma_{\bar{X}}^2=\dfrac{\sigma^2}{n}"

"100=\dfrac{200}{2}"